OpenStudy (anonymous):

(vertex form of a parabola) Write equations in vertex form with the given information. a. vertex of (7,9), a=1 b. vertical stretch of 2, vertex of (-1,-3) c. reflected across the x axis, vertex of (5,0)

OpenStudy (anonymous):

so vertex form is y=a(x-h)^2+k a is the vertical stretch factor if a is negative it gets reflected across the x axis (h,k) is the vertex can you try plugging some things in?

OpenStudy (anonymous):

so it would be \[1(x-7)^{2}2+9\] ???

OpenStudy (anonymous):

was the extra 2 a typo? and since its just a 1 you can just not even write it y=(x-7)^2+9

OpenStudy (anonymous):

What do you mean extra 2 doesnt ^ mean squared?

OpenStudy (anonymous):

Yes, but you have another 2 after the parentheses

OpenStudy (anonymous):

well you wrote y=a(x-h)^2+k is there not suppose to be a 2 at all?

OpenStudy (anonymous):

yes there is but just an exponent this is vertex form \[y=a(x-h)^{2}+k\] plugging things in from number 1 should give you \[y=(x-7)^{2}+9\] not \[y=(x-7)^{2}2+9\]

OpenStudy (anonymous):

ok thats the answer to a i understand that but what about b and c?

OpenStudy (anonymous):

Try it out, what do you think b is?

OpenStudy (anonymous):

do we do the same as the first one? so instead of 1 plug in 2? so 2(x+1)^-3 ??

OpenStudy (anonymous):

really close \[y=2(x+1)^{2}-3\]

OpenStudy (anonymous):

isnt that what i wrote (not to sound rude) and then do you solve it from there or leave it as it is?

OpenStudy (anonymous):

Question asks for vertex form and its in that form. you had the -3 as the exponent, 2 is the exponent remember vertex form is \[ y=a(x-h)^{2}+k\] k is the y coordinate of the vertex and it isnt in the exponent

OpenStudy (anonymous):

Oh i understand what i did wrong and do i just leave the answer as it is or do i try to go further

OpenStudy (anonymous):

Nope dont have to go any further. so now try c

OpenStudy (anonymous):

ok so (5,0) would it be (0,-5)

OpenStudy (anonymous):

I have no idea what you just did =p do it just like the others, except to reflect it across the x axis you make 'a' negative

OpenStudy (anonymous):

Since theres no vertical stretch factor just slap a negative sign where a is

OpenStudy (anonymous):

ok so it is\[-(x-5)^{2}+0\]

OpenStudy (anonymous):

Yep and since its just a 0 you can pretend its not even there. \[y=-(x-5)^{2}\]

OpenStudy (anonymous):

Alright thank u so much

OpenStudy (anonymous):

You're welcome :)